OpenStudy (anonymous):

solve the equation and show work -16 + 2/3 x= -8

OpenStudy (anonymous):

-16 +2/3x = -8 2/3x = -8 +16 2/3x = 8 x = 8 * (3/2) x = 24/2 x = 12

OpenStudy (anonymous):

thanks for breaking it down for me I understand it better this way....thanks a lot

OpenStudy (anonymous):

u r welcome. now pls try to solve some on your, post yr answer and check with us if it is correct :))

OpenStudy (anonymous):

ok